40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12050     Entry time: Mon Mar 28 08:30:09 2016
Author: Steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

Recent  EQ 4.8 mag San Felipe, Mexico trips PRM sus damping.

PRM damping restored. PMC locked.

Attachment 1: shaking.png  622 kB  | Hide | Hide all
shaking.png
ELOG V3.1.3-